OpenStudy (anonymous):

I need help with dividing algebraic fractions... Can someone help me? Please?! >_< (a+3)^2/a-3 / 5/(a^2-9)

OpenStudy (ranga):

\[\Large \frac{ \frac{ (a+3)^2 }{ (a-3) } }{ \frac{ 5 }{ a^2-9 } }\] Is this how the problem appears?

OpenStudy (anonymous):

I believe so but for more clarity I'll just put the picture here. Click on the picture to see it clearer.

OpenStudy (ranga):

Completely different problem.

OpenStudy (anonymous):

Sorry, wrong picture!

OpenStudy (ranga):

\[\Large \frac{ (a+3)^2 }{ (a-3) } \div \frac{ 5 }{ (a^2-9) } = \frac{ (a+3)^2 }{ (a-3) } \times \frac{ (a^2-9) }{ 5 }\]

OpenStudy (ranga):

Dividing by a fraction is same as multiplying by its reciprocal. You can factor (a^2 - 9) as (a + 3)(a - 3)

OpenStudy (ranga):

\[\Large \frac{ (a+3)^2 }{ (a-3) } \times \frac{ (a + 3)(a - 3) }{ 5 } = \frac{ (a+3)^3 }{ 5 }\]

OpenStudy (ranga):

^^^ the (a-3) in the denominator cancels out the (a-3) in the numerator. The (a+3)^2 times (a+3) becomes (a+3)^3

OpenStudy (anonymous):

Okay, thank you!